# High-Field Microscale NMR Spectroscopy with NV Centers in Dipolarly-Coupled Samples

### v1

@@ III Results
We assume the sample starting in a completely mixed state. A triggering RF pulse sets the desired initial state $\rho(0)$, a thermal state oriented along the axis perpendicular to $A$ and $B$. From there, the evolution of the nuclear density matrix is simulated using a master equation that includes $T_{2}^{*}=0.2$ s relaxation,
@@
leading to a signal computed as
@@
where $\sigma_{h}=5.2\times 10^{28}\mathrm{m^{-3}}$ is the number density of hydrogen spins for ethanol, and $g\approx~{}4.1$ is a geometric factor that relates the sample magnetization and the magnetic field in the NV location, see Glenn18 ; Munuera23 for further details.
@@
with $\Omega_{\text{NV}}(t)$ the control Rabi frequency, and $C(t)\frac{\sigma_{z}}{2}$ describing potential RF-induced crosstalk on the NVs. Following the protocol devised in this work, they interact with the signals produced by the nuclear spin rotations around axis $A$, $\bar{A}$ and $\bar{B}$ and accumulate a phase determined by the state of the sample magnetization. During the time that corresponds to the delivery of the RF field over the sample, we tranform the accumulated phase into a population difference with a $\pi/2$ pulse and simulate the measurement, after which the sensor is reinitilized and we simulate the detection of the signal produced by the next LG4 block. Finally, a Fourier transform of the measurements provides the spectra displayed in Fig. (III), which proves the ability of our protocol to access the chemical shifts of the molecule.
Figure (III) shows three different experiments with increasing RF intensities. As expected, stronger RF drivings lead to clearer spectra, less distorted by spurious peaks. In particular we simulate RF drivings with Rabi frequencies of $(2\pi)\times$100 kHz, $(2\pi)\times$150 kHz, and $(2\pi)\times$200 kHz, all attainable values by state of the art antennas Herb20 ; Yudilevich23 , and set the Rabi frequency of the MW control at 20 MHz in all cases. As intended, our method leads to resonance peaks centred in $\delta_{i}^{*}$ from which one can extract the target nuclear shifts $\delta_{i}$ using Eq. (5). For comparison, we include the resutls of a fourth simulation using AERIS Munuera23 , which does not incorporate any dipolar coupling suppression technique. In this case, the obtained spectra (black-curve) is distorted as a consequence of the strong nuclear dipolar couplings.
@@ Appendix A Nuclear spin dynamics under RF drivings
where ±$\pm\delta_{j}/\sqrt{3}$ are the parallel components of the shifts with respect to the effective rotation axis P, and its sign is the same as the sign of $\Delta$. Note that any combination of $\Omega$ and $\Delta$ that complies with the Lee-Goldburg condition produces the described decoupling effect. In particular, for a given intensity of the RF field, this can be detuned from the top and from the bottom with respect to the Larmor. Moreover, the previous derivation is valid for any phase $\alpha$ of the RF field. This freedom has been exploited to develop more elaborated control schemes that concatenate various RF fields, such as the LG4 sequence implemented in our protocol.
@@
Finally, substituting the expression for each axis operator of Eq. (II.1), obtain the propagator
@@
From this expression, we reach the final effective Hamiltonian (4) after rearranging the terms
@@ Appendix B Accumulated Phase
As stated in the main text, the signals received by the NV adhere to a general form Eq.(3). When a two pulse CPMG sequence is applied on the NV sensor (see Fig.(1)) with $\pi$ pulses applied at times $t_{1}$ and $t_{2}$, the phase accumulated by the NV at stage $k$ is:
[…]
@@ Appendix C Analytical expression
Substituting in Eq. (8), we obtain
@@
where $D_{\varphi}=\frac{-\gamma_{e}\hbar^{2}\gamma_{h}^{2}\mu_{0}g\sin{\left(\theta_{\rm LG}\right)}}{8\bar{\Omega}\pi^{2}k_{B}T}\sqrt{\left(\frac{\cos{\varphi}\sin{\alpha}}{\sqrt{3}}-\cos{\alpha}\sin{\varphi}\right)^{2}+\frac{\left(\sqrt{3}\cos{\alpha}\cos{\varphi}+\sin{\alpha}\sin{\varphi}\right)^{2}}{2+\cos{\left(2\alpha\right)}}}$, $\rho_{i}$ is the spin density of the $i$th nucleus, and $\beta_{\varphi}=\arctan{\frac{3\left(\sqrt{3}\cos{\alpha}\cos{\varphi}+\sin{\alpha}\sin{\varphi}\right)}{\sqrt{2+\cos{(2\alpha)}}\left(\sqrt{3}\cos{\varphi}\sin{\alpha}-3\cos{\alpha}\sin{\varphi}\right)}}$. Here, $g$ is a geometric factor that relates the sample geometry with the signal amplitude in the NV site, $k_{B}$ is the Boltzmann constant, and $T$ is the temperature. See Glenn18 ; Munuera23 for further details on the signal amplitude expression. It can be checked that $\varphi_{\rm opt}$ does indeed maximize $D_{\varphi}$. The total accumulated phase of the three drivings $A$, $\bar{A}$, $B$ is simply $3\Phi_{k}$, provided that $\nu_{\varphi}=\mu-\beta_{\varphi}$ is the same in the three stages, which in our case we choose to add up to 00.

### v2

@@ Results
We assume the sample starting in a completely mixed state. A triggering RF pulse sets the desired initial state $\rho(0)$, a thermal state oriented along the axis perpendicular to $A$ and $B$. From there, the evolution of the nuclear density matrix is simulated via the master equation
@@
where $T^{*}_{2}=0.2$ s is chosen to approximate the intrinsic coherence time in the absence of dipolar couplings (which are explicitly included in the Hamiltonian). This equation leads to a signal computed as
@@
where $\sigma_{h}=5.2\times 10^{28}\mathrm{m^{-3}}$ is the number density of hydrogen spins for ethanol, and $g\approx\leavevmode\nobreak\ 4.1$ is a geometric factor that relates the sample magnetization and the magnetic field in the NV location, see [1, 2] for further details.
@@
with $\Omega_{\text{NV}}(t)$ the control Rabi frequency, and $C(t)\frac{\sigma_{z}}{2}$ describing potential RF-induced crosstalk on the NVs. Following the protocol devised in this work, they interact with the signals produced by the nuclear spin rotations around axis $A$, $\bar{A}$ and $\bar{B}$ and accumulate a phase determined by the state of the sample magnetization. During the time that corresponds to the delivery of the $B$ RF field over the sample, we transform the accumulated phase into a population difference with a $\pi/2$ pulse and simulate the measurement, after which the sensor is reinitialized and the protocol repeats. Finally, a Fourier transform of the measurements provides the spectra displayed in Fig. 4, which proves the ability of our protocol to access the chemical shifts of the molecule.
Figure 4 shows three different experiments with increasing RF intensities. As expected, stronger RF drivings lead to clearer spectra, less distorted by spurious peaks which arise due to incomplete averaging of dipolar couplings. In particular we simulate RF drivings with Rabi frequencies of $(2\pi)\times$100 kHz, $(2\pi)\times$150 kHz, and $(2\pi)\times$200 kHz, all attainable values by state of the art antennas [24, 25], and set the Rabi frequency of the MW control at 20 MHz in all cases. Note that when the RF Rabi frequency exceeds the 17 KHz dipolar coupling strength by an order of magnitude, the resulting spectra become significantly cleaner and more resolved. As intended, our method leads to resonance peaks centred in $\delta_{i}^{*}$ from which one can extract the target nuclear shifts $\delta_{i}$ using Eq. (5). For comparison, we include the results of a fourth simulation using AERIS [2], which does not incorporate any dipolar coupling suppression technique. In this case, the obtained spectra (black-curve) is distorted as a consequence of the strong nuclear dipolar couplings.
@@ Acknowledgements
Data availability
The code used to produce the results in this study is available at https://github.com/carlosmunueraj/NV-LG4-DD-samples.
@@ Appendix A Nuclear spin dynamics under RF drivings
where $\pm\delta_{j}/\sqrt{3}$ are the parallel components of the shifts with respect to the effective rotation axis P, and its sign is the same as the sign of $\Delta$. Note that any combination of $\Omega$ and $\Delta$ that complies with the Lee-Goldburg condition produces the described decoupling effect. In particular, for a given intensity of the RF field, this can be detuned from the top and from the bottom with respect to the Larmor. Moreover, the previous derivation is valid for any phase $\alpha$ of the RF field. This freedom has been exploited to develop more elaborated control schemes that concatenate various RF fields, such as the LG4 sequence implemented in our protocol.
@@
Finally, substituting the expression for each axis operator of Eq. (2) in the main text, we obtain the propagator
@@
From this expression, we reach the final effective Hamiltonian after rearranging the terms
@@ Appendix B Accumulated Phase
As stated in the main text, the signals received by the NV adhere to a general form Eq. (3) in the main text. When a two pulse CPMG sequence is applied on the NV sensor (see Fig. (1) in the main text) with $\pi$ pulses applied at times $t_{1}$ and $t_{2}$, the phase accumulated by the NV at stage $k$ is:
$\Phi_{k}=\int_{0}^{t_{1}}\left[|\gamma_{e}|\Gamma_{k}\cos{\left(\bar{\Omega}t+\phi_{k}\right)}+b_{k}\right]dt-\int_{t_{1}}^{t_{2}}\left[|\gamma_{e}|\Gamma_{k}\cos{\left(\bar{\Omega}t+\phi_{k}\right)}+b_{k}\right]dt+\int_{t_{2}}^{T}\left[|\gamma_{e}|\Gamma_{k}\cos{\left(\bar{\Omega}t+\phi_{k}\right)}+b_{k}\right]dt,$ (24)
@@ Appendix C Analytical expression
Substituting in Eq. (8) of the main text, we obtain
@@
where $D_{\varphi}=\frac{-\gamma_{e}\hbar^{2}\gamma_{h}^{2}\mu_{0}g\sin{\left(\theta_{\rm LG}\right)}}{8\bar{\Omega}\pi^{2}k_{B}T}\sqrt{\left(\frac{\cos{\varphi}\sin{\alpha}}{\sqrt{3}}-\cos{\alpha}\sin{\varphi}\right)^{2}+\frac{\left(\sqrt{3}\cos{\alpha}\cos{\varphi}+\sin{\alpha}\sin{\varphi}\right)^{2}}{2+\cos{\left(2\alpha\right)}}}$, $\rho_{i}$ is the spin density of the $i$th nucleus, and $\beta_{\varphi}=\arctan{\frac{3\left(\sqrt{3}\cos{\alpha}\cos{\varphi}+\sin{\alpha}\sin{\varphi}\right)}{\sqrt{2+\cos{(2\alpha)}}\left(\sqrt{3}\cos{\varphi}\sin{\alpha}-3\cos{\alpha}\sin{\varphi}\right)}}$. Here, $g$ is a geometric factor that relates the sample geometry with the signal amplitude in the NV site, $k_{B}$ is the Boltzmann constant, and $T$ is the temperature. See [1, 2] for further details on the signal amplitude expression. It can be checked that $\varphi_{\rm opt}$ does indeed maximize $D_{\varphi}$. The total accumulated phase of the three drivings $A$, $\bar{A}$, $B$ is simply $3\Phi_{k}$, provided that $\nu_{\varphi}=\mu-\beta_{\varphi}$ is the same in the three stages, which in our case we choose to add up to 0.